# Semantic Segmentation Algorithm Based on Light Field and LiDAR Fusion

### v1

@@ Abstract
Semantic segmentation serves as a cornerstone of scene understanding in autonomous driving but continues to face significant challenges under complex conditions such as occlusion. Light field and LiDAR modalities provide complementary visual and spatial cues that are beneficial for robust perception; however, their effective integration is hindered by limited viewpoint diversity and inherent modality discrepancies. To address these challenges, the first multimodal semantic segmentation dataset integrating light field data and point cloud data is proposed. Based on this dataset, we proposed a multi-modal light field point-cloud fusion segmentation network(Mlpfseg), incorporating feature completion and depth perception to segment both camera images and LiDAR point clouds simultaneously. The feature completion module addresses the density mismatch between point clouds and image pixels by performing differential reconstruction of point-cloud feature maps, enhancing the fusion of these modalities. The depth perception module improves the segmentation of occluded objects by reinforcing attention scores for better occlusion awareness. Our method outperforms image-only segmentation by 1.71 Mean Intersection over Union(mIoU) and point cloud-only segmentation by 2.38 mIoU, demonstrating its effectiveness.
@@ I Introduction
As a fundamental task in computer vision, semantic segmentation is crucial for a wide range of applications, including autonomous driving [1], road detection [2], and medical image processing [3]. Existing semantic segmentation methods can be divided into image-based semantic segmentation[4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] and LiDAR-point-cloud-based semantic segmentation[18, 19, 20, 21, 22, 23, 24, 25] according to different types of input data. Image-based semantic segmentation aims to assign a specific semantic label to each pixel in the image, while LiDAR-point-cloud-based semantic segmentation aims to assign a specific semantic label to each point in the point cloud.
Images provide rich color and texture information, but lack accurate 3D spatial structure and are highly sensitive to lighting conditions. In contrast, LiDAR point clouds offer precise spatial geometric data and are less affected by lighting due to their reliance on wavelengths but lack color and texture information. Therefore, effectively fusing images and point clouds is crucial to leverage the strengths of both modalities, enabling more accurate and reliable semantic segmentation.
Recent works, such as 2DPASS [26] and Mseg3D [27], have fused LiDAR point clouds with images to improve 3D segmentation accuracy. CMNeXt [28] integrates LiDAR point cloud benefits for 2D image segmentation. These fusion methods have advanced both 2D and 3D semantic segmentation. However, despite using multiple modalities, they ultimately segment based on a single modality, failing to fully exploit both. When LiDAR point clouds are fused with images but only the images are segmented, the disparity in density between them reduces fusion effectiveness. Sparse point clouds often hinder image segmentation, rather than helping. Moreover, although current multimodal datasets include both cameras and LiDAR, the sparsity of LiDAR point clouds provides limited supplementary information for occluded objects. Additionally, the multiple cameras in existing multimodal datasets often lack overlapping fields of view, making it difficult to effectively complement occluded regions.
Light field images, with multiple viewpoints and significant overlap, provide enhanced occlusion perception by capturing more angular information. By recording different visible parts of occluded objects from various perspectives, which provide more comprehensive cues for occluded content. For example, UrbanLF [4] provides multiple sub-aperture images for richer angular data. However, existing datasets captured using camera arrays offer a large baseline, which is beneficial for exploiting angular information in outdoor scenes, they typically lack semantic annotations, making them unsuitable for segmentation tasks. In contrast, datasets collected using light field cameras provide annotated data, yet their small baseline limits angular diversity and reduces the effectiveness of multi-view fusion. Moreover, most of these datasets only annotate the central view, further restricting methods that aim to leverage full-view light field information.
@@
To address the issues identified above, we constructed TrafficScene [29], the first dataset with semantic annotations that includes both light field images and LiDAR point cloud data. Unlike previous datasets, all viewpoints of the light field are annotated, enabling effective information supplementation for occluded and small objects through multi-view consistency. To integrate light field data and point cloud data effectively, we propose a novel light filed and point cloud fusion segmentation algorithm that aims to fully leverage the complementary strengths of both modalities through simultaneous segmentation of the light filed and point cloud data. To overcome the challenge of poor light filed image segmentation caused by the varying sparsity between point cloud and light filed image during fusion, a pixel-point feature fusion interpolation module is proposed. This module interpolates the features of point clouds projected onto the light filed image plane and subsequently fuses them, thus mitigating the negative impact of sparse point clouds on light filed image segmentation. To enhance the recognition of occluded objects, we introduce a depth difference perception module, which leverages depth information to perceive occlusions.
@@
1. TrafficScene, the first multimodal dataset for semantic segmentation that incorporates light field modalities. Captured using a unique 3×3 camera array with a 30 cm baseline, TrafficScene provides comprehensive semantic annotations across all light field viewpoints, enabling effective multi-view information utilization.
2. We propose the first simultaneous light filed and point cloud segmentation method, Multimodal Light Filed Point Cloud Fusion Segmentation Method (Mlpfseg), it enhances the full integration of point clouds and images and improves the perception of occluded objects through the pixel-point feature fusion interpolation module and the depth difference perception module. Compared to single-image semantic segmentation, the mIoU improves by 1.71. Compared to light field image segmentation, the mIoU improves by 2.37, and compared to multimodal 3D semantic segmentation, the mIoU improves by 2.38.
@@ II Related Works
In this section, we introduce existing light field semantic segmentation datasets and semantic segmentation methods. In this section, we introduce existing semantic segmentation datasets, image semantic segmentation methods, light field semantic segmentation methods, point cloud semantic segmentation methods and multimodal fusion semantic segmentation methods.
@@ II-A Semantic Segmentation Dataset
Among existing semantic segmentation datasets, image-based datasets[30, 31, 32] typically rely on a single perspective, which limits their ability to capture occluded or small objects effectively. Light field image semantic segmentation datasets[4] leverage multiple viewpoints to partially address occlusions; however, they are usually collected using light field cameras with narrow baselines between views, and only the central view is annotated with semantic labels. This limits their effectiveness in supplementing occluded and small object information. Point cloud semantic segmentation datasets[33, 34] provide accurate 3D spatial information and are effective in recognizing occluded objects, but they are inherently sparse, making it difficult to achieve high segmentation accuracy. Existing multimodal semantic segmentation datasets[35, 36, 37, 38, 39] utilize cross-modal fusion to improve recognition of small and occluded objects to some extent. However, since they also rely on a single viewpoint, their ability to comprehensively supplement occluded or small object information remains limited.
@@ II-B Image Semantic Segmentation Methods
Early semantic segmentation algorithms relied on manual feature extraction [6], but with deep learning advancements, recent methods use deep learning for feature extraction. The Fully Convolutional Network (FCN)[7] pioneered deep learning in image segmentation with an end-to-end network using convolutional downsampling and bilinear interpolation. The Pyramid Scene Parsing Network (PSPNet) [8] introduced pyramid pooling for varied object sizes, enhancing precision across scales. DeepLabV3 [9] improved multi-scale object segmentation with various sampling rates for multi-scale contextual information. Recent transformer-based approaches have furthered accuracy in image semantic segmentation. OCRNet [10] employed attention mechanisms for enhanced precision, and Mask2former [11] used masked attention for precise segmentation of smaller objects. SegFormer [12] integrated convolution with MLP for improved object connectivity. Despite theseadvances, single-image semantic segmentation struggles in occlusion or color similarity scenarios due to its limited spatial geometric information capture.
@@ II-C Light Field Semantic Segmentation Methods
Recent algorithms in light field semantic segmentation utilized spatial and angular data to enhance precision, especially in occlusion detail. Chen et al. [5] used CNNs with an angular model and ASPP in light field image segmentation. Sheng et al. [4] stacked images to utilize information from different viewpoints, while Cong et al. [13] applied attention mechanisms and depth maps for enhanced feature extraction. Zhang et al. [14] integrated various perspectives using feature rectification and fusion modules. [15] proposes LF-IENet++, a light field semantic segmentation network that effectively handles multi-baseline disparities via feature integration and propagation. Despite these advancements, the effectiveness of these methods is constrained by the small baseline of light field cameras, limiting the additional information from different viewpoints.
@@ II-D Point Cloud Semantic Segmentation Methods
Unlike images, semantic segmentation of point clouds aims to assign specific categories to each point cloud. Point clouds have gained various uses in autonomous driving due to their unique and accurate three-dimensional spatial structure. There are three main approaches to semantic segmentation of point clouds: directly operating on points, segment after projection, and based on voxels. Methods that directly operate on points, such as PointNet [18] and PointNet++ [19], utilize multi-layer perceptrons to extract features from the point clouds, combining local and global features to classify each point. JSNet++[25] boosts 3D point cloud segmentation via dynamic convolutions and spatial-channel correlation modeling. Projection-based methods map point cloud data onto 2D image coordinates via projection, then apply classic 2D convolutional network architectures for segmentation. Typical projection methods include spherical projection [20] and bird’s-eye projection [21]. A notable algorithm in this category is SqueezeSeg [22], which first converts point clouds into front views using spherical projection before performing segmentation. Recently, voxel-based methods have gained attention. These methods divide the 3D world into voxels and use 3D convolution operations within each voxel block, ultimately obtaining segmentation results for each point in each voxel through upsampling. MinkowskiNet [23] introduced sparse convolution, which efficiently processes high-dimensional sparse data, while SPVCNN [24] divides voxels into regular blocks and applies sparse convolution operations. Methods focused solely on point clouds for 3D semantic segmentation face challenges due to their inherent sparsity. This sparsity leads to a lack of information about small and occluded objects, resulting in a lower segmentation accuracy for these objects.
@@ II-E Multimodal Fusion Semantic Segmentation Methods
Recent algorithms combine images and point clouds to boost segmentation accuracy, generally falling into two categories: those that fuse both modalities but segment only the image, and those that fuse both and segment only the point cloud. In image segmentation, methods like CMNeXt [28] project the point cloud onto the image plane for feature fusion, but the sparsity of point cloud data limits effectiveness. For point cloud segmentation, there are two fusion approaches: data-level and feature-level. In data-level fusion, methods like FuseSeg [40] project the RGB image onto the point cloud’s spherical projection and segment the point cloud, though this can lose intrinsic data structure. Feature-level fusion, such as PMF [41] and 2DPASS [26], extracts multi-scale features from both modalities and fuses them. MSeg3D [27] advances fusion by introducing cross-modal attention, yielding the best results in multimodal fusion methods.
Although recent multimodal fusion semantic segmentation approaches leverage multiple modalities for feature fusion, they typically produce segmentation results for a single modality. This limitation hinders the effective integration of complementary information from both images and point clouds, thereby constraining segmentation performance, particularly in the presence of occluded and small-scale objects. Although multimodal fusion outperforms single-modality point cloud approaches in these scenarios, the improvement remains limited due to the inherent perspective constraint of single-camera imaging, which provides only partial scene information. Therefore, the fusion of light field images and LiDAR point clouds can further enhance the segmentation accuracy of occluded and small objects. To address the challenges of insufficient fusion caused by the difference in spatial density between light field images and point clouds, as well as the limited capability of existing networks to accurately segment occluded and small objects, we propose a novel methods: Mlpfseg. Specifically, we introduce the Point-Pixel Feature Fusion Module (PFFM) to effectively integrate sparse point cloud data with dense light field image. In addition, the Depth Difference Perception Module (DDPM) is designed to enhance the recognition of occluded objects by leveraging depth inconsistency cues.
(a) The light field image of the camera
(b) Annotations of the light field image
(c) The point cloud data
(d) The RGB projection of the image
(e) Annotation of the point cloud
Figure 1: Examples of the data we collected.
(a) Collectionvehicle
(b) Collectionequipment
Figure 2: Multimodal acquisition system.
@@ III Proposed Method
In this paper, we introduce the first multimodal dataset incorporating both light field images and LiDAR point clouds. The light field images are captured using a camera array with a large baseline, providing multiple perspectives that complement each other for better occlusion handling. Annotations are provided for all viewpoints to facilitate full integration. Based on this dataset, we propose a fusion algorithm: Multimodal Light Filed Point Cloud Fusion Segmentation Method (Mlpfseg). Through Point-Pixel Feature Fusion Module(PFFM) and Depth Difference Perception Module(DDPM), we enhance the fusion process, significantly improving segmentation accuracy, especially for occluded objects. These advancements will be discussed in the following sections.
@@ III-A Multimodal TrafficScene Dataset
TrafficScene represents the first multimodal dataset that offers jointly annotated point clouds and full-viewpoint light field images. The following will introduce the collection, calibration, annotation, and statistical analysis of the multimodal dataset. A multimodal dataset of 5607 light field images and 623 frame point clouds was constructed in traffic scenarios using a 3×3 FLIR BFS-PGE 16S2C camera array[42] with a 30 cm baseline and CH128X1 LiDAR[43]. Each frame of the point cloud has more than 60,000 points. The multimodal data acquisition system is shown in Fig. 2. This setup enhances angular diversity and depth perception, aiding in the detection of small or occluded objects and improving semantic segmentation accuracy in complex outdoor scenes. Compared to conventional light field systems, it reduces rain artifacts and improves edge detection. Accurate spatial alignment between modalities is ensured via pairwise calibration. Our aerial perspective captures varied scenes during peak traffic hours, yielding high-resolution (1440×1080) multimodal data. Additional details regarding the dataset are provided in the supplementary material. The dataset is available for download at: https://pan.baidu.com/s/1pY9uY0JP52IeJXmgvfeVcQ.
All light field viewpoints are distortion-corrected and annotated with 15 semantic categories using CVAT[44] As shown in Fig. 3. Annotations are projected onto the LiDAR data and manually refined, making this the first multimodal dataset with aligned semantic labels across both modalities. Our dataset encompasses five representative traffic scenarios—parking lots, urban roads, vegetated roads, roadside areas adjacent to buildings, and roads with obstacles—as well as five types of common traffic participants, including cyclists, pedestrians, buses, cars, and bicycles. Background classes like vegetation and roads dominate, while rich annotations of traffic-specific elements enhance urban scene understanding, offering significant potential for autonomous driving and intelligent transportation systems.
@@ III-B Semantic Segmentation Algorithm Based on Multimodal Data Fusion
For the light field image branch, the input consists of light field images $\{\mathcal{L}_{1},\mathcal{L}_{2},\cdots,\mathcal{L}_{n}\}$, where each image has the size of $\mathbb{R}^{3\times H\times W}$ and $n$ denotes the number of input camera viewpoints. To extract the image features, we employ a weight-shared HRNet-48[45], which enables efficient multi-scale feature representation and enhances the model’s ability to capture detailed spatial information, obtaining the corresponding viewpoint features $\{F_{img1},F_{img2},\cdots,F_{imgn}\}$. The size of each viewpoint feature is $\mathbb{R}^{c_{img}\times h\times w}$, here $c$ represents the number of feature channels; $h$ denotes the height of the feature map; $w$ represents the width of the feature map.
For the point cloud branch, the input point cloud is denoted as $P_{point}\in\mathbb{R}^{N\times 4}$, where $N$ represents the number of points in the point cloud, and 4 corresponds to $\{x_{i},y_{i},z_{i},r_{i}\}$, which represent the 3D spatial coordinates and the refractive index of the $i$-th point, respectively. The point cloud is divided into voxels, and the coordinates of the $i$-th point assigned to the $k$-th voxel are given by $Voxel_{k}=\left\{\left(\left\lfloor\frac{x_{i}}{r_{l}}\right\rfloor,\left\lfloor\frac{y_{i}}{r_{l}}\right\rfloor,\left\lfloor\frac{z_{i}}{r_{l}}\right\rfloor\right)\right\}\in\mathbb{R}^{N\times 3}$. Here, sparse convolution is utilized to extract voxel features through SPVCNN[24], resulting in $F_{voxel}^{l}\in R^{N_{1}\times c_{p}}$, where $F_{voxel}^{l}$ represents the features of the voxels extracted at the $l$-th layer, $N_{1}$ denotes the number of non-empty voxels, and $c_{p}$ represents the number of feature channels per voxel. After obtaining the voxel features at the $l$-th layer, the corresponding features for each point can be derived through interpolation:
$F_{point}^{l}=\sum_{i=1}^{3}\hat{w}_{i}\cdot F_{voxel}^{l}\quad,$ (1)
where $\hat{w}_{i}=\frac{w_{i}}{\sum_{j=1}^{3}w_{j}}$ and $w_{i}=\frac{1}{d(p,v_{i})+\epsilon}$. Here, $p$ represents the point to be interpolated; $v_{i}$ denotes the index of the nearest neighboring voxel; $d(p,v_{i})$ represents the distance between the point and the center of the voxel. The normalized weight is denoted by $\hat{w}_{i}$. The interpolated point features are represented as $F_{point}^{l}\in\mathbb{R}^{N\times c_{p}}$, where $N$ is the number of points; $c_{p}$ is the number of feature channels per point. A small constant $\epsilon=1\times 10^{-8}$ is added to prevent division by zero.
@@
On this basis, a Depth Difference Perception Module (DDPM) is proposed, with the input predicted depth map for each image $D_{pred}$ and the sparse depth map $D_{sparse}$ presenting the depth values for 3D LiDAR coordinates projected onto the image plane. By utilizing depth difference perception, we obtain the attention score map $\hat{D}_{diff}$ for the occluded objects and send it into the PFFM to optimize the representation in the sparse point cloud module. The detailed description of DDPM will be presented in subsection Depth Difference Perception Module (DDPM).
Ultimately, by inputting the $F_{voxel}^{l}$ for each layer , Mseg3D [27], including multi-scale feature extraction modules and context information fusion modules, and the segmentation head are applied to obtain the fused output $\hat{y}_{img}$ in the image branch and the output $\hat{y}_{point}$ in the point cloud branch.

### v2

# Geometry-Aware Cross Modal Alignment for Light Field–LiDAR Semantic Segmentation

@@ Abstract
Semantic segmentation serves as a cornerstone of scene understanding in complex real-world environments. Multimodal sensing, particularly the fusion of light field and LiDAR data, provides complementary appearance and geometric cues, but it remains challenging due to cross-modal discrepancies and limited viewpoint consistency. To address these challenges, a multimodal semantic segmentation dataset integrating light field data and point cloud data with annotations is presented. Based on this dataset, this paper introduces a multimodal light field–LiDAR fusion network (Mlpfseg) for joint semantic segmentation of images and point clouds. The proposed framework introduces a feature completion module to alleviate the density mismatch between sparse point clouds and dense image representations via differential feature reconstruction, and a depth perception module to enhance occlusion-aware feature learning through attention refinement. Experimental results demonstrate that the proposed method achieves 92.38 mean Intersection over Union (mIoU) on point cloud segmentation and 84.97 mIoU on image segmentation, outperforming existing multimodal methods and improving over the baseline by 2.38 mIoU and 3.75 mIoU, respectively.
@@ I Introduction
As a fundamental task in computer vision, semantic segmentation plays a key role in applications such as autonomous driving [1], road detection [2], and medical image analysis [3]. Existing approaches have evolved from traditional machine learning methods to deep learning-based single-modal and multimodal paradigms.
Early methods relied on handcrafted features and classical machine learning techniques such as clustering and support vector machines. With the success of deep learning, single-modal approaches have become dominant, operating on different sensing modalities, including RGB images [4], LiDAR point clouds [5] and infrared images [6]. Image-based methods focus on pixel-wise labeling, while LiDAR-based methods predict semantic labels for 3D points.
RGB images provide rich appearance cues but lack 3D geometry and are sensitive to illumination changes, while LiDAR point clouds offer accurate spatial structure but suffer from sparsity and missing texture information. These complementary characteristics have motivated extensive research on multimodal semantic segmentation for improved robustness. However, existing fusion methods such as 2DPASS [7] and Mseg3D [8] often rely on asymmetric or weak cross-modal interaction, limiting full exploitation of complementary information. In addition, LiDAR sparsity may even degrade image segmentation performance due to modality imbalance, rather than consistently providing benefits. More importantly, conventional multimodal datasets based on multi-camera systems typically lack sufficient overlapping views, making it difficult to recover occluded regions in complex traffic scenes.
To address this limitation, light field imaging has recently attracted increasing attention. Unlike conventional imaging systems, light field data provides dense multi-view observations with strong viewpoint overlap, enabling more complete perception of occluded objects. UrbanLF [9] provides rich sub-aperture images but lacks semantic annotations for segmentation. Other light field datasets include annotations but are often constrained by limited baseline and angular resolution, with supervision typically restricted to central views, preventing full exploitation of multi-view information.
@@
To address the issues identified above, we constructed TrafficScene [10], the first dataset with semantic annotations that includes both light field images and LiDAR point cloud data. Unlike previous datasets, all viewpoints of the light field are annotated, enabling effective information supplementation for occluded and small objects through multi-view consistency. To effectively integrate light field and point cloud data, we propose a novel fusion-based semantic segmentation framework that jointly performs segmentation on both modalities, thereby better exploiting their complementary characteristics. To alleviate the degradation of light field image segmentation caused by the sparsity discrepancy between point clouds and images, we design a pixel–point feature fusion interpolation module. This module interpolates the features of point clouds projected onto the light field image plane and subsequently fuses them, thus mitigating the negative impact of sparse point clouds on light field image segmentation. To enhance the recognition of occluded objects, we introduce a depth difference perception module, which leverages depth information to perceive occlusions.
@@
1. TrafficScene, the first multimodal dataset for semantic segmentation that incorporates light field images and LiDAR point clouds with annotations. Captured using a unique 3×3 camera array with a LiDAR sensor, TrafficScene provides comprehensive semantic annotations across all light field viewpoints, enabling effective multi-view information utilization.
2. We propose a novel light field and LiDAR fusion framework, Multimodal Light field Point Cloud Fusion Segmentation Method (Mlpfseg). It enhances the full integration of point clouds and images and improves the perception of occluded objects through the Point-Pixel Feature Fusion Module and the Depth Difference Perception Module.
3. Extensive experiments demonstrate that Mlpfseg consistently outperforms representative baseline and state-of-the-art methods for single-image, light field image, and multimodal semantic segmentation on the TrafficScene dataset. In particular, it achieves improvements of +3.75 and +2.38 mIoU over the baseline under different settings, while also surpassing previous state-of-the-art methods by up to +1.9 mIoU. Additional results on two benchmark datasets further validate the robustness and generalization capability of the proposed method.
@@ II Related Works
In this section, we introduce existing semantic segmentation datasets, image semantic segmentation methods, light field semantic segmentation methods, point cloud semantic segmentation methods and multimodal fusion semantic segmentation methods.
@@ II-A Semantic Segmentation Dataset
Image-based datasets[11, 12, 13] typically rely on a single perspective, limiting their ability to capture complete scene information. Light field datasets[9] exploit multi-view information to alleviate occlusion issues; however, their narrow baselines and limited annotation (usually only the central view) restrict their effectiveness. Point cloud datasets[14, 15] provide accurate 3D spatial structure but suffer from sparsity. Multimodal datasets[16, 17, 18, 19, 20] improve performance through cross-modal fusion, yet their reliance on single-view imaging still constrains comprehensive scene understanding.
@@ II-B Image Semantic Segmentation Methods
Early approaches relied on handcrafted features [21], while modern methods are dominated by deep learning. Representative CNN-based methods, such as FCN [4], PSPNet [22], and DeepLabV3 [23], improve segmentation performance through multi-scale context modeling. More recently, transformer-based approaches, including OCRNet [24], Mask2Former [25], and SegFormer [26], further enhance global context modeling and representation capability. Despite these advances, single-image methods remain limited in handling occlusions and geometrically ambiguous regions due to the lack of explicit 3D structural information.
@@ II-C Light Field Semantic Segmentation Methods
Light field semantic segmentation leverages both spatial and angular information to improve pixel-level recognition, particularly in occluded or geometrically ambiguous regions. Existing methods exploit multi-view consistency and angular cues to enhance feature representation. For example, Chen et al. [27] employed CNNs with angular modeling and ASPP for multi-scale context extraction, while Sheng et al. [9] aggregated sub-aperture images to capture complementary viewpoints. Subsequent works incorporated attention mechanisms and depth information [28], as well as feature rectification and cross-view fusion [29], to improve consistency across views. More recently, LF-IENet++ [30] introduced an effective feature integration strategy to address multi-baseline disparities. However, these methods are fundamentally limited by the narrow baseline of light field cameras, which restricts the available geometric and depth cues.
@@ II-D Point Cloud Semantic Segmentation Methods
Point cloud semantic segmentation aims to assign semantic labels to individual 3D points, leveraging precise geometric structure for scene understanding. Existing approaches can be broadly categorized into point-based, projection-based, and voxel-based methods. Point-based methods, such as PointNet [31] and PointNet++ [32], directly learn features from raw points using multilayer perceptrons. Projection-based methods map point clouds onto 2D representations and apply 2D CNNs [33, 34]. Voxel-based methods discretize space into volumetric grids and employ sparse 3D convolutions, as demonstrated by MinkowskiNet [35] and SPVCNN [36]. Despite their effectiveness, point cloud methods suffer from inherent sparsity and irregularity, leading to degraded performance in small or occluded object regions.
@@ II-E Multimodal Fusion Semantic Segmentation Methods
Multimodal semantic segmentation integrates complementary information from images and point clouds to improve performance. Existing methods can be roughly divided into image-oriented and point cloud-oriented approaches. Image-oriented methods, such as CMNeXt [29], project point clouds onto the image plane for feature enhancement, but are limited by the sparsity of projected points. Point cloud-oriented methods adopt either data-level fusion, e.g., FuseSeg [37], or feature-level fusion, such as PMF [38] and 2DPASS [7]. More advanced approaches, such as MSeg3D [8], introduce cross-modal attention to improve fusion effectiveness. Nevertheless, most existing methods produce segmentation results for only a single modality, limiting the full exploitation of cross-modal information. Moreover, the reliance on single-view images restricts comprehensive scene understanding.
((a))
((b))
((c))
((d))
((e))
Figure 1: Multimodal data samples from our dataset. (a) Light field images captured by the cameras. (b) Pixel-level annotations of the light field images. (c) LiDAR point cloud data. (d) RGB projection of the point cloud onto the image plane. (e) Corresponding annotations of the point cloud data.
Figure 2: Multimodal data acquisition system.
@@ III Proposed Method
Building upon our previously established multimodal dataset TrafficScene[10] that integrates light field images and LiDAR point clouds, this paper investigates semantic segmentation in complex outdoor environments. The light field modality provides rich multi-view cues that are effective for handling occlusions, while LiDAR preserves precise 3D geometric structure. Based on this representation, we propose a unified fusion framework, termed Multimodal Light Field Point Cloud Fusion Segmentation (Mlpfseg). The framework consists of two key components: the Point–Pixel Feature Fusion Module (PFFM), which enables fine-grained cross-modal interaction, and the Depth Difference Perception Module (DDPM), which explicitly models cross-modal depth inconsistencies. These designs jointly improve segmentation performance, particularly in occluded regions. The remainder of this section details the proposed framework.
@@ III-A Multimodal TrafficScene Dataset
TrafficScene is, to the best of our knowledge, the first multimodal dataset that provides jointly annotated LiDAR point clouds and full-view light field images for semantic segmentation. It was originally introduced in our previous work [10], and this paper presents an extended version with improved annotation quality, enhanced cross-modal alignment, and increased diversity of traffic scenarios.
The dataset contains 5607 light field images and 623 LiDAR frames collected using a multimodal acquisition system [39] in real-world traffic environments, as shown in Fig. 2. The system consists of a $\times3$ FLIR BFS-PGE 16S2C camera array [40] with a 30 cm baseline and a CH128X1 LiDAR sensor [41], enabling accurate synchronization and calibration across modalities. The dataset is available for download at: https://github.com/yxJiang0125/TrafficScene-Dataset
All light field views are distortion-corrected and annotated with 15 semantic categories using CVAT [42], including pedestrians, vehicles, cyclists, traffic signs, roads, sidewalks, vegetation, and other urban elements (Fig. 3). Cross-view consistency is enforced via label propagation based on feature correspondence and geometric constraints. LiDAR annotations are obtained through calibrated projection and further refined using the Xtreme1 platform [43]. For overlapping projections, semantic labels are merged via majority voting to improve robustness under occlusion.
Overall, the dataset provides synchronized multimodal annotations for both light field images and LiDAR point clouds, enabling comprehensive 2D–3D semantic segmentation. Multi-view light field imagery alleviates LiDAR sparsity, while LiDAR provides complementary geometric structure, forming a strong benchmark for cross-modal scene understanding.
@@ III-B Semantic Segmentation Algorithm Based on Multimodal Data Fusion
For the light field image branch, the input consists of light field images $\{\mathcal{L}_{1},\mathcal{L}_{2},\cdots,\mathcal{L}_{n}\}$, where each image has size $\mathbb{R}^{3\times H\times W}$ and $n$ denotes the number of camera viewpoints. We employ a weight-shared HRNet-48 [44] for multi-scale feature extraction, producing viewpoint features $\{F_{\text{img}1},F_{\text{img}2},\cdots,F_{\text{img}n}\}\in\mathbb{R}^{c_{\text{img}}\times h\times w}$.
For the point cloud branch, the input point cloud is denoted as $P_{\text{point}}\in\mathbb{R}^{N\times 4}$, where $N$ is the number of points and each point is represented by $\{x_{i},y_{i},z_{i},r_{i}\}$. The point cloud is voxelized by discretizing coordinates into voxel grids:
$Voxel_{k}=\left\{\left(\left\lfloor\frac{x_{i}}{r_{l}}\right\rfloor,\left\lfloor\frac{y_{i}}{r_{l}}\right\rfloor,\left\lfloor\frac{z_{i}}{r_{l}}\right\rfloor\right)\right\}\in\mathbb{R}^{N\times 3}.$
Voxel features are extracted using SPVCNN [36], yielding $F_{\text{voxel}}^{l}\in\mathbb{R}^{N_{1}\times c_{p}}$, where $N_{1}$ denotes the number of non-empty voxels. Point-level features are obtained via interpolation:
$F_{\text{point}}^{l}=\sum_{i=1}^{3}\hat{w}_{i}\cdot F_{\text{voxel}}^{l},$ (1)
where $\hat{w}_{i}=\frac{w_{i}}{\sum_{j=1}^{3}w_{j}}$ and $w_{i}=\frac{1}{d(p,v_{i})+\epsilon}$. Here $d(p,v_{i})$ denotes the distance between point $p$ and its neighboring voxel center $v_{i}$, and $\epsilon=10^{-8}$ prevents numerical instability.
@@
On this basis, a Depth Difference Perception Module (DDPM) is proposed, with the input predicted depth map for each image $D_{pred}$ and the sparse depth map $D_{sparse}$ presenting the depth values for 3D LiDAR coordinates projected onto the image plane. By utilizing depth difference perception, we obtain the attention score map $\hat{D}_{diff}$ for the occluded objects and send it into the PFFM to optimize the representation in the sparse point cloud module. The detailed description of DDPM will be presented in subsection Depth Difference Perception Module (DDPM). Ultimately, by inputting the $F_{voxel}^{l}$ for each layer, Mseg3D [8], including multi-scale feature extraction modules and context information fusion modules, and the segmentation head are applied to obtain the fused output $\hat{y}_{img}$ in the image branch and the output $\hat{y}_{point}$ in the point cloud branch.